OpenStudy (anonymous):

you throw a water balloon vertically upwards with a velocity of 6 m/s. how long until the balloon comes back and breaks in your face? can someone please help me!

OpenStudy (anonymous):

helpppp

OpenStudy (anonymous):

vel with which its thrown up, u=6m/s when it returns to the same point from which its thrown its velocity will be same in magnitude but opposite in direction... that is velocity with which it will splash to your face is, v=-6m/s we have v=u+at here a= -9.8m/s^2 so -6=6+(-9.8*t) t=12/9.8